### v1

@@ 4.1 Synthethic Detection Track
[…]
ACC / AUC Grag2021 Corvi2022 Wu2023 Corvi2022* Ours*
[…]
Discussion – On the Practical test set, our model outperforms the Wu2023 approach, yielding almost 7% improvement in ACC and 4% in AUC on average. Importantly, our approach exhibits strong generalization capabilities, performing effectively on unseen generators (DreamBooth, NightCafe, and YiJian) even without explicit training data from them and outperforming Wu2023 on two of them while preserving the same results on the third. The Grag2021 method, trained only on StyleGAN images, underperforms on YiJian (community model, part of NightCafe [1]). It achieves lower accuracy than 50%, suggesting the detector’s inability to generalize over the different generators’ examples, giving predictions similar to a random classifier. The Corvi2022 approach almost matches the performance of Wu2023 on Stable Diffusion examples. The other Corvi2022 model trained on the ImagiNet training set has a great improvement in the result compared to previous state-of-the-art detectors, signifying the quality of the dataset and the importance of data diversity. However, this detector has limited generalization abilities to some of the unseen generators (NightCafe, YiJian) when compared to our baseline.
Our baseline performance remains consistent on Corvi’s benchmark as there is no statistically significant difference to the Corvi2022 model – mean ACC almost 87% and mean AUC at $0.95$. It shows a substantial improvement of 12% in ACC on DALL·E 2 examples since it is trained on DALL·E 3 images. The results on StyleGAN3 and StyleGAN2 are increased by 1-2%. In comparison to our model, Grag2021 has a higher AUC on almost all GAN sections of the benchmark as it is trained only on GAN examples. However, its performance decreases significantly on all other generators. The Corvi2022 model trained on ImagiNet performs almost the same as our baseline and Corvi2022 detector.
[…]
ACC / AUC Grag2021 Corvi2022 Wu2023 Corvi2022* Ours*
[…]
On the ImagiNet benchmark, the SelfCon approach achieves the highest average ACC and AUC across all detectors. It is important to highlight that while the Corvi2022 model, trained on our ImagiNet dataset, demonstrates improved performance compared to other models, our training method further enhances this performance and stands out in terms of generalizability. This is evident in the near-random classifier accuracy exhibited by other models on proprietary generative models such as Midjourney and DALL·E 3, highlighting their limited ability to generalize beyond the benchmarks lacking data diversity. Further analysis of performance across different content types and specific generators can be found in the supplementary material.
We recognize the critical importance of inference speed in real-world applications. While our proposed model may not achieve the absolute fastest inference (Table 6), it significantly outperforms the state-of-the-art Corvi2022 model, reducing its inference time by half while maintaining comparable accuracy on its test set and surpassing it on others. It signifies the ability of our baseline to balance performance and efficiency, making it a more practical solution for time-sensitive applications.
Table 6: Models inference time in milliseconds for an image with size $448\times 448$ on a single NVIDIA RTX 4090 GPU.
@@
4.2 Model Identification Track
@@
Discussion – Interestingly, even aggressive perturbations such as resize and compression do not significantly harm the performance of the model identifier (Table 7). The results confirm that each generator has its unique "fingerprints", which are studied in other works [12]. The model identification track is open for evaluation of other novel detectors. Further analysis of the model robustness under perturbations is placed in the supplementary materials.
Table 7: Results of our detector on ImagiNet model identification track in terms of accuracy. The AUC is $0.9980$
GAN SD Midjourney DALL·E 3 Mean
0.9968 0.9683 0.9228 0.9160 0.9510
5 Conclusion
In conclusion, we introduce ImagiNet, a high-resolution balanced dataset designed to address the limitations of previous sets in the field of synthetic image detection. They often lack diverse content types and multiple generators, leading to biased detectors. By involving a wide range of content types and generators, including proprietary models, ImagiNet enables the training of more generalizable and robust detectors. The quality of our ImagiNet dataset is demonstrated by the fact that a model trained on it, even with a previous state-of-the-art method, surpasses current detectors on several benchmarks. In addition, we utilize a self-contrastive learning approach that further enhances the model’s generalizability, achieving state-of-the-art results under challenging social network conditions, such as image resizing and compression. Notably, our model reaches this performance while maintaining fast inference speed, making it a practical solution for real-world applications.
6 Acknowledgements
We would like to thank Sangmin Bae for the advice provided. We also acknowledge the America for Bulgaria and Beautiful Science foundations for partially funding the computational resources used as a part of the Diffground OSS project. The Google ML Developer Programs team supported this work by providing Google Cloud and Colab credits.
Appendix 0.A Assessment of Potential Biases
To investigate the influence of specific content types and synthetic image generators on the synthetic image detector and identify potential biases, we conducted an ablation study inspired by Leave-One-Out Cross-Validation (LOOCV). Separate models were trained, each with one category (content type, generator) excluded from its training data, while maintaining equal training data overall. The isolation of the specific category influence allowed us to identify potential biases through drastic changes in performance when tested on the unseen group of examples.
0.A.1 Content Type
From the synthetic images in our ImagiNet dataset, we focused on those generated by Stable Diffusion due to its presence in all image subcategories, thus eliminating potential generator-specific biases. We sampled a balanced subset containing 4500 real and 4500 synthetic (Stable Diffusion only) images per subcategory (photos, paintings, faces). For each model, we used a ResNet-18 architecture, training it from scratch for 200 epochs to avoid any biases from pre-trained models. Each model was trained on 18000 images with one category left out. This lets us analyze how each type of content affects the detector’s performance by comparing it to a baseline model trained on all categories. For evaluation, we sample 1000 real and 1000 synthetic images for each category. [^3] We investigated a potential bias toward the resolution of real images across different content type groups. However, our analysis revealed no significant bias.
Figure 4: Real and synthetic accuracy on the different models that are trained by leaving one content type out and the baseline trained on all content types.
Figure 5: Mean accuracy and AUC on the different models trained by leaving one content type out and the baseline trained on all content types.
The plots presented in Figure 4 and Figure 5 demonstrate that models trained by excluding a specific content type exhibit overfitting and generally lower synthetic accuracy when tested on that content type. Notably, the "Except Faces" model overfits the real image distribution, suggesting that bias is introduced not only by synthetic images but also by real images. The AUC plot in Figure 5 reveals high variance from expected values for the "Except Painting" and "Except Faces" models on their respective content types, highlighting the inability to distinguish between the real and synthetic classes at all possible thresholds. This suggests that training on diverse content types is essential for mitigating bias. The baseline model, trained on all types, does not overfit on the test set and reduces the bias, indicating the suitable setting for these most common content types found on the World Wide Web.
0.A.2 Synthetic Image Generator
To study biases introduced by different synthetic image generators while excluding biases from varying content types, we focus exclusively on the photos category from our ImagiNet dataset. We sampled a balanced subset containing 4,500 real photographs and 4,500 synthetic photos for each generator (StyleGAN-XL, ProGAN, SD v2.1/SDXL v1.0). We trained three individual ResNet-18 detector models from scratch for 200 epochs, each on 18,000 images, excluding one generator’s examples from the training (respectively detectors’ names are ’Except StyleGAN-XL’, ’Except ProGAN’, and ’Except SD’). This allows us to assess the performance of each detector in the absence of influence from a specific generator through the training data, and compare it to a baseline model trained on all generators. For evaluation, we sampled 1,000 real and 1,000 synthetic images per generator.
Figure 6: Real and synthetic accuracy on the different models that are trained by leaving one synthetic generator out and the baseline trained on images from all synthetic generators.
Figure 7: Mean accuracy and AUC on the different models trained by leaving one synthetic generator out and the baseline trained on examples from all generators.
As shown in the "Real ACC" plot of Figure 6, the diversity of real photos used during training does not significantly affect the accuracy of synthetic image detection, as expected, since we are measuring bias towards synthetic data. In contrast, in the "Synthetic ACC" plot of Figure 6, models tested on images from unseen generators perform poorly, with some barely reaching 0.04% - 0.09% accuracy. This highlights that models struggle to generalize to synthetic images from unseen generators during training. Even models trained on GAN examples, like ’Except StyleGAN-XL’ (trained on ProGAN and Stable Diffusion), cannot generalize to images from StyleGAN-XL. The AUC plot in Figure 7 further demonstrates that detectors trained without images from a specific generator fail to distinguish real from the excluded synthetic source, even at various thresholds, indicating the presence of bias related to the synthetic image generator. The performance of the baseline model, trained on diverse synthetic sources, suggests that this diversity enhances the overall detector’s generalization.
Appendix 0.B Performance of Detectors by Content Type and Specific Generator
Following the procedure outlined in the synthetic track, we conducted an evaluation of detector bias on ImagiNet, focusing on both accuracy and AUC as key performance metrics and grouping the images by content type and specific generator. All detectors included in the comparison are Grag2021 [16], Corvi2022 [12], Wu2023 [52], Corvi2022* [12] trained on our ImagiNet dataset and our SelfCon method. For each content type and specific generator, our model achieved the highest performance in both accuracy and AUC when compared to others (Table 8 and Table 9).
To further investigate potential biases, we examined the range of accuracy and AUC values (difference between maximum and minimum for all categories) across different content types and generators for each model. We chose range as our metric because it clearly shows how much performance varies across different conditions. A smaller range means the model performs more consistently, while a larger range suggests potential biases affecting specific types of content or generators. Our model consistently showed the smallest range in both accuracy and AUC, even compared to Corvi2022 trained on the same data. This indicates that our method is more generalizable and less susceptible to these types of bias.
[…]
ACC / AUC Grag2021 Corvi2022 Wu2023 Corvi2022* Ours*
[…]
ACC / AUC Grag2021 Corvi2022 Wu2023 Corvi2022* Ours*
[…]
Appendix 0.C Model Identification Under Perturbations
@@
Appendix 0.D Visualizations via Relational Dimensionality Reduction
High-dimensional data often poses challenges for analysis since relationships between data points can’t be visualized and interpreted directly. To address this, we employ dimensionality reduction techniques. Specifically, we use an autoencoder architecture [5] to project the data into a lower-dimensional space, allowing us to potentially identify decision boundaries. The model consists of two parts: an Encoder $\mathrm{g}_{enc}()$ and a Decoder $\mathrm{g}_{dec}()$. Each layer in the network is linear with ReLU activation. Batched vectors $R_{in}\in\mathbb{R}^{b\times n}$ where $b$ is the batch size - are inputted to the model. The dimensionality $n$ decreases layer-wise until it reaches dimension $h$ dimensionality of encoded vectors $R_{enc}\in\mathbb{R}^{h}$. In our case, $h=2$ since we want to plot the vectors as points in 2D space. Following that, the decoder increases the dimensionality, outputting reconstructed vector $R_{out}\in\mathbb{R}^{n}$. To achieve accurately encoded vectors, we need to minimize the Mean Squared Error (MSE) between $R_{in}$ and $R_{out}$. Due to the inability of reconstruction error to preserve the relations between vectors in the high-dimensional space, we introduce relational loss. This objective preserves the pairwise distances between vectors in the batch after decoding and also in lower-dimensional mapping. Our method adopts a similar framework to Meng et al. [33]. Instead of calculating the MSE between Gram matrices, we focus directly on the pairwise $l_{2}$ distances between original input vectors and their decoded counterparts, as this preserves the magnitude of the distances and allows us to maintain the original positional relationships in the data. We then calculate the MSE over these pairwise distances, the value of which indicates the average squared error in reconstructing the individual relationships present within the original data. Hence, the objective is the following:
@@ Appendix 0.D Visualizations via Relational Dimensionality Reduction
$\begin{split}\mathcal{L}_{rdra}=(1-\alpha)\frac{\sum_{i=1}^{b}\sum_{j=1}^{b}(R_{in[i,j]}-R_{out[i,j]})^{2}}{b^{2}}+\\ \alpha\frac{\sum_{i=1}^{b}\sum_{j=1}^{b}({\lVert R_{out[i,:]}-R_{out[j,:]}\rVert}^{2}-{\lVert R_{in[i,:]}-R_{in[j,:]}\rVert}^{2})^{2}}{b^{2}}\end{split}$ (5)
@@
Appendix 0.E Image Generation and Prompt Engineering
@@
All the images are generated with different random seeds for the initial noise to achieve diverse generated content. To remove potential bias, during the image generation, we utilize different schedulers for generation [8] – Euler Discrete, Euler Ancestral [22], DPM-Solver [30], PNDM [28].
Appendix 0.F Additional Training Details

### v2

@@ Introduction
GAN 0.6889 / 0.8403 0.6822 / 0.8033 0.6508 / 0.6971 0.8534 / 0.9416 0.9372 / 0.9886
SD 0.5140 / 0.5217 0.6112 / 0.6851 0.6367 / 0.6718 0.8693 / 0.9582 0.9608 / 0.9922
Midjourney 0.4958 / 0.5022 0.5826 / 0.6092 0.5326 / 0.5289 0.8880 / 0.9658 0.9652 / 0.9949
DALL·E 3 0.4128 / 0.3905 0.5180 / 0.5270 0.5368 / 0.5482 0.8906 / 0.9759 0.9724 / 0.9963
Mean 0.5279 / 0.5637 0.5985 / 0.6562 0.5892 / 0.6115 0.8753 / 0.9604 0.9589 / 0.9930
Table 3: ImagiNet test set evaluation – best ACC/AUC. * means trained on ImagiNet.
ACC / AUC Wu2023 Ours*
DreamBooth 0.9049 / 0.9733 0.9601 / 0.9950
MidjoruneyV4 0.8907 / 0.9495 0.9675 / 0.9959
MidjourneyV5 0.8540 / 0.9224 0.9745 / 0.9991
NightCafe 0.8962 / 0.9652 0.8931 / 0.9644
StableAI 0.8806 / 0.9534 0.9574 / 0.9947
YiJian 0.8392 / 0.9233 0.9045 / 0.9726
Mean 0.8776 / 0.9479 0.9428 / 0.9870
Table 4: Practical test set (Wu, Zhou, and Zhang 2023) evaluation – best ACC/AUC. * means trained on ImagiNet.
ACC / AUC Corvi2022 Corvi2022* Ours*
ProGAN 0.9117 / 0.9994 0.9030 / 0.9995 0.8974 / 0.9991
StyleGAN2 0.8662 / 0.9455 0.8675 / 0.9479 0.8884 / 0.9759
StyleGAN3 0.8557 / 0.9416 0.8705 / 0.9440 0.8824 / 0.9707
BigGAN 0.8952 / 0.9699 0.8980 / 0.9882 0.8934 / 0.9864
EG3D 0.9062 / 0.9756 0.8450 / 0.9160 0.8964 / 0.9913
Taming Tran 0.9112 / 0.9902 0.8538 / 0.9278 0.8829 / 0.9651
DALL·E Mini 0.9117 / 0.9914 0.9015 / 0.9792 0.8924 / 0.9786
DALL·E 2 0.6507 / 0.7590 0.7370 / 0.8302 0.7729 / 0.8590
GLIDE 0.9062 / 0.9780 0.8730 / 0.9429 0.8539 / 0.9347
Latent Diff 0.9117 / 0.9998 0.9017 / 0.9989 0.8959 / 0.9902
Stable Diff 0.9117 / 0.9999 0.9030 / 0.9998 0.8969 / 0.9956
ADM 0.7927 / 0.8772 0.7875 / 0.8710 0.7704 / 0.8550
Mean 0.8692 / 0.9523 0.8618 / 0.9446 0.8686 / 0.9585
Table 5: Corvi test set (Corvi et al. 2022) evaluation – best ACC/AUC. * means trained on ImagiNet.
First, we evaluate the described baseline against existing synthetic datasets. Then, we examine the importance of balancing content types in ImagiNet for the performance of detectors.
Baseline – During the first stage, the backbone is optimized with SGD (Ruder 2017) for 400 epochs with batch size $N=200$ on the ImagiNet training set. The initial learning rate of $0.005$ is warmed up linearly (Ma and Yarats 2021) for 10 epochs and is cosine annealed (Loshchilov and Hutter 2017) afterwards. The second stage continues for 5 epochs with AdamW optimizer (Loshchilov and Hutter 2017) and constant learning rate $0.0001$, weight decay $0.001$, $\beta_{1}=0.9$ and $\beta_{2}=0.99$. After the pre-training procedure, we visualize the model representations of the images in the test set by applying Autoencoder dimensionality reduction (Meng et al. 2017). The plots in Figure 2 show the ability of our model to cluster each generator’s images.
As shown in Table 3, the baseline achieves AUC of up to 0.99 and balanced accuracy over 95% on ImagiNet. To demonstrate its generalization abilities we evaluate zero-shot performance on the datasets from (Wu, Zhou, and Zhang 2023) in Table 4, and (Corvi et al. 2022) in Table  5. Our baseline is able to outperform the original method of Wu2023 and remains comparable on Corvi2022’s benchmark. The baseline shows a substantial improvement of 12% in ACC on DALL·E 2 examples since it is trained on DALL·E 3 images. The results on StyleGAN3 and StyleGAN2 are increased by 1-2%. Table 6 presents a comparison of the inference time of our detector with previous models. We also train the model proposed in Corvi2022 on ImagiNet to demonstrate that the balanced dataset elicits generalizable performance regardless of the architecture and training procedure.
Content Type Balancing – To investigate the influence of specific content types and identify potential biases, we conducted an ablation study inspired by Leave-One-Out Cross-Validation (LOOCV). Separate models were trained, each with one content type excluded from its training data, while maintaining equal training data overall. The isolation of the specific category influence allows us to identify potential biases through drastic changes in performance when tested on the unseen group of examples.
From the synthetic images in our ImagiNet dataset, we focused on those generated by Stable Diffusion due to its presence in all image subcategories, thus eliminating potential generator-specific biases. We sampled a balanced subset containing 4500 real and 4500 synthetic (Stable Diffusion only) images per subcategory (photos, paintings, faces). For each model, we used a ResNet-18 architecture, training it from scratch for 200 epochs to avoid any biases from pre-trained models. Each model was trained on 18000 images with one category left out. For evaluation, we sample 1000 real and 1000 synthetic images for each category. [^2] Our analysis revealed no significant bias toward the resolution of real images across different content type groups.
Figure 3 demonstrate that models trained by excluding a specific content type exhibit overfitting and generally lower synthetic accuracy when tested on that content type. Notably, the “Except Faces” model overfits the real image distribution, suggesting that bias is introduced not only by synthetic images but also by real images. The AUC plot in Figure 3 reveals high variance from expected values for the “Except Painting” and “Except Faces” models on their respective content types, highlighting the inability to distinguish between the real and synthetic classes at all possible thresholds. This suggests that training on diverse content types is essential for mitigating bias. The baseline model, trained on all types, does not overfit on the test set.
@@
Table 6: Inference time in milliseconds for $448\times 448$ image on RTX 4090 GPU.
Figure 3: Mean accuracy and AUC on the different models trained by leaving one content type out.
Conclusion
In this work: (1) we demonstrate the importance of balancing content types in synthetic image datasets; (2) we provide a modest-in-size but high quality benchmark for training and evaluating synthetic detectors; (3) we provide a strong baseline which generalizes on third-party datasets.
Acknowledgements
We would like to thank Sangmin Bae for the advice provided. We also acknowledge the America for Bulgaria and Beautiful Science foundations for partially funding the computational resources used as a part of the Diffground OSS project. The Google ML Developer Programs team supported this work by providing Google Cloud and Colab credits. This work was partially supported by Bulgarian National Program ”Education with Science”.
Appendix A Model Identification Track
@@
Discussion – Interestingly, even aggressive perturbations such as resize and compression do not significantly harm the performance of the model identifier (Table 7). The results confirm that each generator has its unique ”fingerprints”, which are studied in other works (Corvi et al. 2022). The model identification track is open for evaluation of other novel detectors.
Model Identification Under Perturbations
@@
Table 7: Results of our detector on ImagiNet model identification track - ACC. The AUC is $0.9980$
GAN SD Midjourney DALL·E 3 Mean
0.9968 0.9683 0.9228 0.9160 0.9510
Appendix B Visualizations via Relational Dimensionality Reduction
High-dimensional data often poses challenges for analysis since relationships between data points can’t be visualized and interpreted directly. To address this, we employ dimensionality reduction techniques. Specifically, we use an autoencoder architecture (Bank, Koenigstein, and Giryes 2021) to project the data into a lower-dimensional space, allowing us to potentially identify decision boundaries. The model consists of two parts: an Encoder $\mathrm{g}_{enc}()$ and a Decoder $\mathrm{g}_{dec}()$. Each layer in the network is linear with ReLU activation. Batched vectors $R_{in}\in\mathbb{R}^{b\times n}$ where $b$ is the batch size - are inputted to the model. The dimensionality $n$ decreases layer-wise until it reaches dimension $h$ dimensionality of encoded vectors $R_{enc}\in\mathbb{R}^{h}$. In our case, $h=2$ since we want to plot the vectors as points in 2D space. Following that, the decoder increases the dimensionality, outputting reconstructed vector $R_{out}\in\mathbb{R}^{n}$. To achieve accurately encoded vectors, we need to minimize the Mean Squared Error (MSE) between $R_{in}$ and $R_{out}$. Due to the inability of reconstruction error to preserve the relations between vectors in the high-dimensional space, we introduce relational loss. This objective preserves the pairwise distances between vectors in the batch after decoding and also in lower-dimensional mapping. Our method adopts a similar framework to Meng et al. (2017). Instead of calculating the MSE between Gram matrices, we focus directly on the pairwise $l_{2}$ distances between original input vectors and their decoded counterparts, as this preserves the magnitude of the distances and allows us to maintain the original positional relationships in the data. We then calculate the MSE over these pairwise distances, the value of which indicates the average squared error in reconstructing the individual relationships present within the original data. Hence, the objective is the following:
@@ Appendix B Visualizations via Relational Dimensionality Reduction
$\begin{split}\mathcal{L}_{rdra}=\sum_{i=1}^{b}\sum_{j=1}^{b}\frac{(1-\alpha)}{b^{2}}(R_{in[i,j]}-R_{out[i,j]})^{2}+\\ \frac{\alpha}{b^{2}}({\lVert R_{out[i,:]}-R_{out[j,:]}\rVert}^{2}-{\lVert R_{in[i,:]}-R_{in[j,:]}\rVert}^{2})^{2}\end{split}$ (3)
@@
Appendix C Image Generation and Prompt Engineering
@@
All the images are generated with different random seeds for the initial noise to achieve diverse generated content. To remove potential bias, during the image generation, we utilize different schedulers for generation (Chen 2023) – Euler Discrete, Euler Ancestral (Karras et al. 2022), DPM-Solver (Lu et al. 2022), PNDM (Liu et al. 2022).
Appendix D Additional Training Details